# Pushing the Limits of On-Device Streaming ASR: A Compact, High-Accuracy English Model for Low-Latency Inference

### 16 Apr 2026 v1

@@ 1 Introduction
Automatic Speech Recognition has undergone rapid improvements in recent years, with large-scale models such as OpenAI Whisper [1], NVIDIA NeMo models [2], and Qwen3-ASR [3] pushing the boundaries of accuracy on standard English benchmarks. However, the best-performing models—Qwen3-ASR-1.7B at 5.90% WER, Parakeet TDT-0.6B-v3 at 6.32%, and Canary-1B-v2 at 7.15%—are all primarily batch-oriented architectures requiring 2–7 GB of memory and GPU inference. Deploying high-quality ASR in edge scenarios, where compute is limited to CPUs, latency must be minimal, and memory budgets are tight, remains a significant challenge.
The goal of on-device ASR is to bring speech recognition directly to user devices without reliance on cloud infrastructure. Our primary objective was to identify a high-quality English streaming ASR model that operates under the lowest possible resource utilization—minimizing model size, memory consumption, CPU load, and algorithmic latency without sacrificing transcription quality. This translates to four concrete constraints:
@@ 2.2 Parakeet TDT and Conformer Transducer
We group these two models together as both are transducer-based architectures. NVIDIA’s Parakeet TDT-0.6B-v3 [6] uses a Token-and-Duration Transducer architecture. We tested it extensively across 15+ chunking configurations to characterize the relationship between chunk size, context length, and WER on English benchmarks. The Conformer Transducer XL [8] is a larger English-only conformer transducer model also evaluated in chunked mode.
@@ 4 Batch-Mode Comparison
We first evaluated all models in their native batch (offline) mode to establish accuracy upper bounds. Table 3 shows the results.
@@ 6.2 Quantization Methods
All quantization variants use weight-only block quantization: each weight matrix $W$ is partitioned into contiguous blocks of $b$ elements (block size $b=32$ in our experiments), and each block is independently mapped to $n$-bit integers ($n\in\{4,8\}$) with a per-block scale $s$ and zero-point $z$. Activations remain in FP32 at inference time; the quantized weights are consumed by the fused ONNX MatMulNBits operator, which combines linear dequantization with matrix multiplication. The two quantization algorithms we evaluate differ in how $s$ and $z$ are computed:
Round-To-Nearest (RTN) [11].
@@
To explore beyond the initial quantization, k-quant evaluates 20 candidate scale factors, uniformly spaced in a narrow range around the initial RTN scale, each producing different integer assignments $q_{j}$. For each candidate, the optimal $(s^{*},m^{*})$ is recomputed in closed form, and the candidate with the lowest weighted error is kept.
@@ 7.2 Key Findings
    Quantization preserves accuracy remarkably well. The int8 k-quant variant achieves 8.01% average WER, essentially matching the ONNX FP32 baseline (8.03%), with a 48% reduction in model size. Even the int4 k-quant variant at 0.67 GB (73% smaller) reaches 8.20%, only 0.17% above FP32.
@@
    Aggressive compression is viable. At int4 precision with a model size of just 0.67 GB, the WER remains within 0.2% absolute of the ONNX FP32 baseline (WERR = 1.021, i.e., only 2.1% relative degradation), demonstrating that 4-bit quantization introduces minimal degradation for streaming ASR.
@@
[…]
  5. 5.
    ONNX FP32 vs. PyTorch baseline gap. The ONNX FP32 variant shows a 0.75% absolute WER increase over the PyTorch CUDA baseline (8.03% vs. 7.28%). We suspect that this discrepancy is not caused by the ONNX export itself—as the graph is numerically equivalent, but rather by differences in kernel implementations, which may compound through 48 conformer layers.
@@ 9 Discussion
Quantization preserves accuracy remarkably well. On the full evaluation sets, the int4 k-quant variant achieves 8.20% average WER compared to the 8.03% ONNX FP32 baseline, a degradation of only 0.17% absolute despite a 73% reduction in model size (from 2.47 GB to 0.67 GB). The degradation is distributed across most datasets, with the largest increases on AMI (+0.65%) and SPGISpeech (+0.22%), while TED-LIUM actually shows no degradation (4.65% vs. 4.66%). The int8 variant at 1.28 GB essentially matches FP32 accuracy (8.01% vs. 8.03%), confirming that 8-bit quantization is effectively lossless for this architecture. One notable negative result is the ConvInteger/MatMulInteger variant, which degrades to 10.14% WER despite using the same int4 k-quant weights. Unlike standard quantized operators that dequantize inputs back to floating point before computing, ConvInteger and MatMulInteger perform the entire computation in integer arithmetic, accumulating rounding errors through the encoder’s 48 conformer layers of interleaved convolutions and attention.
@@ 9.2 Concurrent Work
We note that concurrent to this work, Mistral released Voxtral Mini [12], a multimodal model with strong ASR capabilities including streaming support. While promising, its model size ($\sim$4 GB) exceeds the sub-1 GB target for our edge deployment scenario. Similarly, Useful Sensors released Moonshine Streaming [13], a lightweight encoder-decoder model with sliding-window attention designed for on-device streaming. Although its smallest variant fits comfortably within our size constraints, its autoregressive design introduces a different latency profile from the transducer-based system studied here, with output latency that can grow with transcript length even when time-to-first-token is low. In addition, as with other sequence-to-sequence ASR models, it may be more susceptible to repetitions or hallucination-like behaviors in challenging conditions. A comprehensive empirical comparison with these models is left for future work.
@@ A.7 Qwen3-ASR — All Configurations
A.8 Nemotron-0.6B — ONNX Quantization Results
@@
A.9 Nemotron-0.6B — ONNX int4 k-quant at 0.16s Delay

### 19 Apr 2026 v2

@@ 1 Introduction
Automatic Speech Recognition has undergone rapid improvements in recent years, with large-scale models such as OpenAI Whisper [1], NVIDIA NeMo models [2], and Qwen3-ASR [3] pushing the boundaries of accuracy on standard English benchmarks. However, the best-performing models—Qwen3-ASR-1.7B at 5.90% WER, Parakeet TDT-0.6B-v3 at 6.32%, and Canary-1B-v2 at 7.15%—are all primarily batch-oriented architectures requiring 2–7 GB of memory and GPU inference. Deploying high-quality ASR in edge scenarios, where compute is limited to CPUs, latency must be minimal, and memory budgets are tight, remains a significant challenge. Prior work has studied strong ASR architectures, streaming inference strategies, and low-bit quantization, but these directions are often evaluated in isolation. This report focuses on their intersection: identifying a strong streaming architecture under a controlled public benchmark suite and compressing it for practical edge deployment.
Our primary objective was to identify a high-quality English streaming ASR model that operates under the lowest possible resource utilization—minimizing model size, memory consumption, CPU load, and algorithmic latency without sacrificing transcription quality—enabling speech recognition directly on user devices without reliance on cloud infrastructure. This translates to four concrete constraints:
@@ 2.2 Parakeet TDT and Conformer Transducer
We group these two models together as both are transducer-based architectures. NVIDIA’s Parakeet TDT-0.6B-v3 [6] uses a Token-and-Duration Transducer architecture with multilingual support, making it a strong candidate for future language expansion if chunked streaming proves viable. We tested it extensively across 15+ chunking configurations to characterize the relationship between chunk size, context length, and WER on English benchmarks. The Conformer Transducer XL [7] is a larger English-only conformer transducer model also evaluated in chunked mode.
@@ 4 Batch-Mode Comparison
We first evaluated all models in batch (offline) mode to establish accuracy upper bounds. All results reported in this section reflect our own evaluation under the controlled setup described in Section 3; reported values may differ from officially published numbers due to differences in hardware, text normalization, or evaluation methodology. Table 3 shows the results.
@@ 5.3 Comparison with Other Streaming/Chunked Models
  • While Parakeet TDT achieves higher RTFx than Nemotron on GPU (see Table 9 in the appendix), it is notably slower on CPU (RTFx 1.38 vs. 2.46). We hypothesize that Parakeet’s larger chunks benefit from GPU parallelism over wide tensors, whereas on CPU, where parallelism is limited, Nemotron’s smaller, cache-efficient chunks incur less per-step compute and better exploit the CPU memory hierarchy.
@@ 6.2 Quantization Methods
Post-training quantization methods can be broadly divided into calibration-based approaches, which use representative input data to guide quantization decisions, and calibration-free approaches, which determine quantization parameters entirely from weight statistics. Calibration-based methods such as AWQ [11] and GPTQ [12] are most commonly studied in autoregressive LLM settings, where representative calibration data is naturally defined in terms of text prompts. Prior ASR work has also explored low-bit quantization-aware training, including 2-bit QAT for Conformer models on LibriSpeech and large-scale internal data [13], though such approaches require retraining. For a stateful streaming transducer operating on audio input, defining representative calibration data and chunk-level execution conditions is less straightforward, so in this work we focus on calibration-free post-training strategies that require no additional training or calibration data.
All quantization variants use weight-only block quantization: each weight matrix $W$ is partitioned into contiguous blocks of $b$ elements (block size $b=32$ in our experiments), and each block is independently mapped to $n$-bit integers ($n\in\{4,8\}$) with a per-block scale $s$ and zero-point $z$. Activations remain in FP32 at inference time; the quantized weights are consumed by the fused ONNX MatMulNBits operator, which combines linear dequantization with matrix multiplication. We evaluate two calibration-free quantization algorithms that differ in how $s$ and $z$ are computed:
Round-To-Nearest (RTN).
@@
To explore beyond the initial quantization, our k-quant evaluates 20 candidate scale factors, uniformly spaced in a narrow range around the initial RTN scale, each producing different integer assignments $q_{j}$. For each candidate, the optimal $(s^{*},m^{*})$ is recomputed in closed form, and the candidate with the lowest weighted error is kept.
@@ 7.2 Key Findings
    Quantization preserves accuracy remarkably well. The int8 k-quant variant achieves 8.01% average WER, essentially matching the ONNX FP32 baseline (8.03%), with a 48% reduction in model size. Even at int4 precision with a model size of just 0.67 GB (73% smaller), the WER degrades by only 0.17% absolute (2.1% relative), demonstrating that aggressive 4-bit compression is viable for streaming ASR.
@@
    ONNX FP32 vs. PyTorch baseline gap. The ONNX FP32 variant shows a 0.75% absolute WER increase over the PyTorch CUDA baseline (8.03% vs. 7.28%). We suspect that this discrepancy is not caused by the ONNX export itself, since the graph is numerically equivalent, but rather by differences in kernel implementations that may compound through 24 conformer layers.
@@ 9 Discussion
Quantization preserves accuracy remarkably well. On the full evaluation sets, the int4 k-quant variant achieves 8.20% average WER compared to the 8.03% ONNX FP32 baseline, a degradation of only 0.17% absolute despite a 73% reduction in model size (from 2.47 GB to 0.67 GB). The degradation is distributed across most datasets, with the largest increases on AMI (+0.65%) and SPGISpeech (+0.22%), while TED-LIUM actually shows no degradation (4.65% vs. 4.66%). The int8 variant at 1.28 GB essentially matches FP32 accuracy (8.01% vs. 8.03%), confirming that 8-bit quantization is effectively lossless for this architecture. One notable negative result is the ConvInteger/MatMulInteger variant, which degrades to 10.14% WER despite using the same int4 k-quant weights. Unlike standard quantized operators that dequantize inputs back to floating point before computing, ConvInteger and MatMulInteger perform the entire computation in integer arithmetic, accumulating rounding errors through the encoder’s 24 conformer layers of interleaved convolutions and attention.
@@ 9.2 Concurrent Work
We note that concurrent to this work, Mistral released Voxtral Mini [14], a multimodal model with strong ASR capabilities including streaming support. While promising, its model size ($\sim$4 GB) exceeds the sub-1 GB target for our edge deployment scenario. Similarly, Useful Sensors released Moonshine v2 [15], a lightweight encoder-decoder model with sliding-window attention designed for on-device streaming. Although its model variants fit comfortably within our size constraints, its autoregressive design introduces a different latency profile from the transducer-based system studied here, with output latency that can grow with transcript length even when time-to-first-token is low. In addition, as with other sequence-to-sequence ASR models, it may be more susceptible to repetitions or hallucination-like behaviors in challenging conditions. A comprehensive empirical comparison with these models is left for future work.
@@ A.7 Qwen3-ASR — All Configurations
A.8 Cross-Model Streaming Comparison (CPU)
Table 14: Cross-model streaming comparison (CPU, batch_size=1). Best streaming or low-latency configuration per model.
Model Size Delay Avg BSF RTFx AMI Earn. Giga LS-c LS-o SPGI TED VoxP.
A.9 Nemotron-0.6B — ONNX Quantization Results
@@ A.8 Cross-Model Streaming Comparison (CPU)
A.10 Nemotron-0.6B — ONNX int4 k-quant at 0.16s Delay